The perimeter of a rectangular
backyard is 6x + 6 yards. If the width is x yards, find a
binomial that represents the length.

I know that the perimeter of a rectangle is p=2L + 2W but I am not sure how to work out this problem.

Found 3 solutions by richwmiller, stanbon, checkley77:
Answer by richwmiller(9143)

you have
p=2(W+L)
but we are using x for w
p=2(x+L)
we are also told that p= 6x+6
6x+6=2(x+L)
6x+6=2x+2L
4x+6=2L
2x+3=L

Answer by stanbon(57424)

The perimeter of a rectangular backyard is 6x + 6 yards.
If the width is x yards, find a binomial that represents the length.
----
Perimeter = 2(W + L)
---
6x+6 = 2(x + L)
Divide both sides by 2 to get:
3x+3 = x+L
---
Length = 2x+3
==================
Cheers,
Stan H.
==================

Answer by checkley77(12569)

2L+2W=PERIMETER.
W=X
2L+2X=6X+6
2L=6X-2X+6
2L=4X+6
L=(4X+6)/2
L=(2X+3) ANS. FOR THE LENGTH.